# How Redundant Is the Transformer Stack in Speech Representation Models?

### v1

@@ 3.3 Mimicking Networks
Is the transformer stack completely redundant? To answer this question, we substitute the transformer stack with mimicking layers. Based on the two blocks we find during the similarity analysis, we first test two mimicking layers, with the intermediate mimicked layer $i$ being the last layer of the first block. We compare these results to only using one mimicking layer, directly learning the final representations, as well as to using just a linear or transformer layer instead, directly trained for classification. These substitutions lead to a parameter reduction of up to 95% in base models and up to 98% in large models and a reduction of the inference time of up to 94% (see Figure 3 and Figure 4-6). Most models retain more than 95% of the performance while achieving these high reductions in size and therefore also inference time. This is a higher reduction, while maintaining comparable performance, than in recent studies on KD of speech representation models [15, 16, 17, 18] and a much higher reduction in size than in pruning studies on speech representation models [5, 12, 13]. Although our method does not bring significant improvements over other state-of-the-art distillation methods, it demonstrates the high redundancy of the transformer stack and highlights the potential for significant reductions in size and computational load.
@@ 4 Conclusion
Appendix A Appendix
@@
Results of simplification of networks using mimicking networks. In wav2vec large (Figure 4) and wavLM (Figure 5) the models keep over 95% of their original performance while reducing the number of parameters by 95-98% and the inference time by up to 91%. In wavLM large (Figure 6) the performance is still above 90% of the original performance while reducing the size by 98% and the inference time by 94%.
Fig. 4: Simplification of transformer stack using mimicking networks for wav2vec2-large. Reduction in inference time (up to 91%) and number of parameters (up to 98%) using mimicking networks, while maintaining over 95% of the original accuracy ( ). We test transformer and linear layers with different dimensions $z$. Inference time is normalized to 1 (inference time of original model).
Fig. 5: Simplification of transformer stack using mimicking networks for wavLM. Reduction in inference time (up to 90%) and number of parameters (up to 95%) using mimicking networks, while maintaining over 95% of the original accuracy ( ). We test transformer and linear layers with different dimensions $z$. Inference time is normalized to 1 (inference time of original model).
Fig. 6: Simplification of transformer stack using mimicking networks for wavLM-large. Reduction in inference time (up to 94%) and number of parameters (up to 98%) using mimicking networks, while almost maintaining 95% of the original accuracy ( ). We test transformer and linear layers with different dimensions $z$. Inference time is normalized to 1 (inference time of original model).

### v2

@@ 3.3 Mimicking Networks
Is the transformer stack completely redundant? To answer this question, we substitute the transformer stack with mimicking layers. Based on the two blocks we find during the similarity analysis, we first test two mimicking layers, with the intermediate mimicked layer $i$ being the last layer of the first block. We compare these results to only using one mimicking layer, directly learning the final representations, as well as to using just a linear or transformer layer instead, directly trained for classification. These substitutions lead to a parameter reduction of up to 95% in base models and up to 98% in large models and a reduction of the inference time of up to 94% (seeFigure 3 and Figure 4-6 in the appendix). Most models retain more than 95% of the performance while achieving these high reductions in size and therefore also inference time. This is a higher reduction, while maintaining comparable performance, than in recent studies on KD of speech representation models [15, 16, 17, 18] and a much higher reduction in size than in pruning studies on speech representation models [5, 12, 13]. Although our method does not bring significant improvements over other state-of-the-art distillation methods, it demonstrates the high redundancy of the transformer stack and highlights the potential for significant reductions in size and computational load.
@@ 4 Conclusion
Appendix A Appendix / supplemental material
We present the data behind Figure 3 (see Table 1) along with experiments of mimicking networks for the wav2vec2-large, wavLM-small and wavLM-large. Note that the L and T respectively denote linear and transformer layers, while $z$ is the hidden dimension of the layer(s).
@@ Appendix A Appendix / supplemental material
Results of simplification of networks using mimicking networks. In wav2vec2-large (Figure 4 and Table 2) and wavLM-small (Figure 5 and Table 3) the models keep over 95% of their original performance while reducing the number of parameters by 95-98% and the inference time by up to 91%. In wavLM-large (Figure 6 and Table 4) the performance is still above 90% of the original performance while reducing the size by 98% and the inference time by 94%.
Network Layer $N$ $z$ Number of Inference time Accuracy
type type layers parameters (normalized)
Original T 12 - 94577571 1 $0.976\pm 0.002$
Mimicker L 1 32 4851331 0.13 $0.919\pm 0.004$
Mimicker L 2 32 4901283 0.14 $0.925\pm 0.004$
Mimicker L 1 768 5984035 0.14 $0.942\pm 0.003$
Mimicker L 2 768 7165219 0.14 $0.938\pm 0.004$
Mimicker L 1 4096 11105827 0.15 $0.935\pm 0.004$
Mimicker L 2 4096 17402147 0.16 $0.934\pm 0.004$
Mimicker T 1 32 7216707 0.14 $0.936\pm 0.004$
Mimicker T 2 32 9632099 0.16 $0.934\pm 0.004$
Mimicker T 1 768 8347939 0.15 $0.921\pm 0.004$
Mimicker T 2 768 11894563 0.16 $0.942\pm 0.003$
Mimicker T 1 4096 13463075 0.16 $0.945\pm 0.003$
Mimicker T 2 4096 22124835 0.18 $0.944\pm 0.003$
Non-mimicker L 1 32 4851331 0.13 $0.916\pm 0.004$
Non-mimicker L 1 768 5984035 0.14 $0.936\pm 0.004$
Non-mimicker L 1 4096 11105827 0.14 $\mathbf{0.945\pm 0.003}$
Non-mimicker T 1 32 7216707 0.14 $0.93\pm 0.004$
Non-mimicker T 1 768 8347939 0.15 $0.931\pm 0.004$
Non-mimicker T 1 4096 13463075 0.16 $0.925\pm 0.004$
Table 1: wav2vec2-small
Network Layer $N$ $z$ Number of Inference time Accuracy
type type layers parameters (normalized)
Original T 24 - 315700387 1 $0.971\pm 0.002$
Mimicker L 1 32 5064835 0.089 $0.921\pm 0.004$
Mimicker L 2 32 5131427 0.091 $0.924\pm 0.004$
Mimicker L 1 768 6574371 0.09 $0.938\pm 0.004$
Mimicker L 2 768 8149027 0.094 $0.94\pm 0.004$
Mimicker L 1 4096 13400099 0.097 $\mathbf{0.945\pm 0.003}$
Mimicker L 2 4096 21793827 0.11 $0.943\pm 0.003$
Mimicker T 1 32 9267267 0.098 $0.927\pm 0.004$
Mimicker T 2 32 13536355 0.11 $0.931\pm 0.004$
Mimicker T 1 768 10775331 0.099 $0.937\pm 0.004$
Mimicker T 2 768 16552483 0.11 $0.938\pm 0.004$
Mimicker T 1 4096 17594403 0.11 $0.94\pm 0.004$
Mimicker T 2 4096 30190627 0.12 $0.937\pm 0.004$
Non-mimicker L 1 32 5064835 0.088 $0.917\pm 0.004$
Non-mimicker L 1 768 6574371 0.09 $0.937\pm 0.004$
Non-mimicker L 1 4096 13400099 0.097 $0.939\pm 0.004$
Non-mimicker T 1 32 9267267 0.098 $0.925\pm 0.004$
Non-mimicker T 1 768 10775331 0.1 $0.936\pm 0.004$
Non-mimicker T 1 4096 17594403 0.11 $0.929\pm 0.004$
Table 2: wav2vec2-large
Fig. 4: wav2vec2-large
Fig. 5: wavLM-small
Fig. 6: wavLM-large
Network Layer $N$ $z$ Number of Inference time Accuracy
type type layers parameters (normalized)
Original T 12 - 94587795 1 $0.972\pm 0.002$
Mimicker L 1 32 4851331 0.097 $0.889\pm 0.005$
Mimicker L 2 32 4901283 0.099 $0.91\pm 0.004$
Mimicker L 1 768 5984035 0.1 $0.938\pm 0.004$
Mimicker L 2 768 7165219 0.1 $0.936\pm 0.004$
Mimicker L 1 4096 11105827 0.1 $\mathbf{0.939\pm 0.004}$
Mimicker L 2 4096 17402147 0.11 $0.932\pm 0.004$
Mimicker T 1 32 7216707 0.1 $0.903\pm 0.004$
Mimicker T 2 32 9632099 0.11 $0.928\pm 0.004$
Mimicker T 1 768 8347939 0.11 $0.928\pm 0.004$
Mimicker T 2 768 11894563 0.12 $0.933\pm 0.004$
Mimicker T 1 4096 13463075 0.11 $0.928\pm 0.004$
Mimicker T 2 4096 22124835 0.13 $0.932\pm 0.004$
Non-mimicker L 1 32 4851331 0.097 $0.885\pm 0.005$
Non-mimicker L 1 768 5984035 0.1 $0.921\pm 0.004$
Non-mimicker L 1 4096 11105827 0.1 $0.924\pm 0.004$
Non-mimicker T 1 32 7216707 0.1 $0.915\pm 0.004$
Non-mimicker T 1 768 8347939 0.11 $0.915\pm 0.004$
Non-mimicker T 1 4096 13463075 0.11 $0.923\pm 0.004$
Table 3: wavLM-small
Network Layer $N$ $z$ Number of Inference time Accuracy
type type layers parameters (normalized)
Original T 24 - 315724515 1 $0.989\pm 0.002$
Mimicker L 1 32 5070979 0.064 $0.859\pm 0.005$
Mimicker L 2 32 5137571 0.065 $0.876\pm 0.005$
Mimicker L 1 768 6580515 0.065 $0.906\pm 0.004$
Mimicker L 2 768 8155171 0.067 $\mathbf{0.93\pm 0.004}$
Mimicker L 1 4096 13406243 0.069 $0.915\pm 0.004$
Mimicker L 2 4096 21799971 0.075 $0.913\pm 0.004$
Mimicker T 1 32 9273411 0.07 $0.849\pm 0.005$
Mimicker T 2 32 13542499 0.076 $0.828\pm 0.006$
Mimicker T 1 768 10781475 0.07 $0.888\pm 0.005$
Mimicker T 2 768 16558627 0.079 $0.829\pm 0.006$
Mimicker T 1 4096 17600547 0.075 $0.91\pm 0.004$
Mimicker T 2 4096 30196771 0.086 $0.821\pm 0.006$
Non-mimicker L 1 32 5070979 0.064 $0.88\pm 0.005$
Non-mimicker L 1 768 6580515 0.064 $0.893\pm 0.005$
Non-mimicker L 1 4096 13406243 0.069 $0.905\pm 0.004$
Non-mimicker T 1 32 9273411 0.07 $0.907\pm 0.004$
Non-mimicker T 1 768 10781475 0.071 $0.907\pm 0.004$
Non-mimicker T 1 4096 17600547 0.075 $0.89\pm 0.005$
Table 4: wavLM-large